# Simple homotopy types of even dimensional manifolds

### v2

@@ 2.2. The Whitehead group
Define the Whitehead group of a group $G$ to be $\operatorname{Wh}(G)=K_{1}(\mathbb{Z}G)/{\pm G}$, where the map ±$\pm G\to K_{1}(\mathbb{Z}G)$ is the composition ±$\pm G\subseteq GL_{1}(\mathbb{Z}G)\subseteq\operatorname{GL}(\mathbb{Z}G)\to K_{1}(\mathbb{Z}G)$.
Note that $\operatorname{Wh}$ is a functor from the category of groups to the category of abelian groups. The map induced by a homomorphism $\theta\colon G\rightarrow H$ will be denoted by $\theta_{*}\colon\operatorname{Wh}(G)\rightarrow\operatorname{Wh}(H)$. Similarly, for a continuous map $f\colon X\rightarrow Y$ between topological spaces, the map induced by $\pi_{1}(f)\colon\pi_{1}(X)\rightarrow\pi_{1}(Y)$ will be denoted by $f_{*}\colon\operatorname{Wh}(\pi_{1}(X))\rightarrow\operatorname{Wh}(\pi_{1}(Y))$.
The isomorphism type of $\operatorname{Wh}(G)$ is known in some cases. For us the following examples will be relevant.
Proposition 2.9 (Stallings [St65]).
If $G$ is a finitely generated free group, then $\operatorname{Wh}(G)=0$.
@@
The Whitehead group $\operatorname{Wh}(G)$ of a group $G$ is equipped with a natural involution, i.e. an automorphism $x\mapsto\overline{x}$ such that $\overline{\overline{x}}=x$. Equivalently, it has a $\mathbb{Z}C_{2}$-module structure, where the generator of $C_{2}$ acts by the involution. We will describe this involution below.
@@
If $G$ is a group and $w\colon G\to\{\pm 1\}$ is an orientation character, i.e. a homomorphism from $G$ to {±$\{\pm 1\}$, then the integral group ring $\mathbb{Z}G$ has an involution given by $\sum_{i=1}^{k}n_{i}g_{i}\mapsto\sum_{i=1}^{k}w(g_{i})n_{i}g_{i}^{-1}$ for $n_{i}\in\mathbb{Z}$ and $g_{i}\in G$. The resulting involution on $K_{1}(\mathbb{Z}G)$ preserves ±$\pm G$ and so induces an involution on $\operatorname{Wh}(G)$.
@@
We will write $\operatorname{Wh}(G,w)$ for the abelian group $\operatorname{Wh}(G)$ equipped with the involution determined by the orientation character $w\colon G\to\{\pm 1\}$. This can equivalently be regarded as a $\mathbb{Z}C_{2}$-module. When $w\equiv 1$ is the trivial homomorphism, we will omit it from the notation and write $\operatorname{Wh}(G)$ for $\operatorname{Wh}(G,1)$.
Note that the choice of $w$ only affects the involution, and $\operatorname{Wh}(G,w)$ is equal to $\operatorname{Wh}(G)$ as an abelian group. Next we define two subgroups of $\operatorname{Wh}(G,w)$ which will play an important rôle in the rest of this article.
Definition 2.12.
For a group $G$ and a homomorphism $w\colon G\to\{\pm 1\}$, define:
$\mathcal{J}_{n}(G,w)$$\coloneqq\{y\in\operatorname{Wh}(G,w)\mid y=-(-1)^{n}\overline{y}\}\leq\operatorname{Wh}(G,w)$
$\mathcal{I}_{n}(G,w)$$\coloneqq\{x-(-1)^{n}\overline{x}\mid x\in\operatorname{Wh}(G,w)\}\leq\operatorname{Wh}(G,w)$
We have $\mathcal{I}_{n}(G,w)\leq\mathcal{J}_{n}(G,w)$, and it follows from the definition of the Tate cohomology groups (see Proposition 9.5) that there is a canonical isomorphism
@@ 2.3. The Whitehead torsion of a chain homotopy equivalence
Now we define the Whitehead torsion in the algebraic setting. Let $G$ be a group and let $f\colon C_{*}\to D_{*}$ be a chain homotopy equivalence between finitely generated, free, based (left) $\mathbb{Z}G$-module chain complexes. Consider the algebraic mapping cone $(\mathscr{C}(f),\partial_{*})$, which is also a finitely generated, free, based $\mathbb{Z}G$-module chain complex. Since $f$ is a chain homotopy equivalence, $\mathscr{C}(f)$ is chain contractible; see e.g. [Ranicki-AGS, Proposition 3.14]. Choose a chain contraction $s\colon\mathscr{C}(f)_{*}\to\mathscr{C}(f)_{*+1}$. Now consider the modules
$\mathscr{C}(f)_{\operatorname{odd}}\coloneqq\bigoplus_{i=0}^{\infty}\mathscr{C}(f)_{2i+1}\quad\text{ and }\quad\mathscr{C}(f)_{\operatorname{even}}\coloneqq\bigoplus_{i=0}^{\infty}\mathscr{C}(f)_{2i}.$
@@
$\tau(f)\coloneqq\Bigl{[}\bigoplus_{i=0}^{\infty}\bigl{(}\partial_{2i+1}+s_{2i+1}\bigr{)}\colon\mathscr{C}(f)_{\operatorname{odd}}\to\mathscr{C}(f)_{\operatorname{even}}\Bigr{]}\in\operatorname{Wh}(G).$
This equivalence class is independent of the choice of the chain contraction $s$ [Co73, (15.3)].
@@
Since $\mathscr{C}(f)_{\operatorname{odd}}$ and $\mathscr{C}(f)_{\operatorname{even}}$ are finitely generated, the sum in the definition of $\tau(f)$ is a finite sum.
The equivalence class $\tau(f)$ remains invariant under permutations of the bases of $C_{*}$ and $D_{*}$, or if a basis element is multiplied by $(-1)$ or an element of $G$ [Co73, (15.2) and (10.3)].
Now we list some useful facts about $\tau$. We fix the group $G$, and all chain complexes will be assumed to be finitely generated, free, based, left $\mathbb{Z}G$-module chain complexes.
@@
The following lemma gives a useful way to compute Whitehead torsion in favourable special cases. We will apply it in Section 6.
Lemma 2.18.
Let $f\colon C_{*}\to D_{*}$ be a chain map such that $f_{i}\colon C_{i}\to D_{i}$ is an isomorphism for each $i$. Then $[f_{i}]\in\operatorname{Wh}(G)$ for each $i$ and $\tau(f)=\sum_{i=0}^{\infty}(-1)^{i}[f_{i}]$.
Proof.
Let $n=\max\left\{i\mid C_{i}\neq 0\right\}$, we will prove the statement by induction on $n$. If $n=0$, then $\mathscr{C}(f)_{\operatorname{odd}}=\mathscr{C}(f)_{1}\cong C_{0}$, $\mathscr{C}(f)_{\operatorname{even}}=\mathscr{C}(f)_{0}\cong D_{0}$, and $\partial_{1}=f_{0}$, so $\tau(f)=[f_{0}]$ by Definition 2.13.
For the induction step define $C^{\prime}_{*}$ by $C^{\prime}_{i}=C_{i}$ if $i<n$ and $C_{i}=0$ otherwise, and define $C^{\prime\prime}_{*}$ by $C^{\prime\prime}_{n}=C_{n}$ and $C_{i}=0$ if $i\neq n$. Define $D^{\prime}_{*}$, $D^{\prime\prime}_{*}$, $f^{\prime}\colon C^{\prime}_{*}\rightarrow D^{\prime}_{*}$, and $f^{\prime\prime}\colon C^{\prime\prime}_{*}\rightarrow D^{\prime\prime}_{*}$ analogously, and apply Lemma 2.17. ∎
For a chain complex $C_{*}$ and an integer $k$ we will denote by $C_{k+*}$ the same chain complex with shifted grading $i\mapsto C_{k+i}$. Similarly, the cochain complex $C_{k-*}$ is defined by changing the grading to $i\mapsto C_{k-i}$. The notation $C^{k+i}$ and $C^{k-*}$ is defined analogously for a cochain complex $C^{*}$.
@@
Let $f\colon C^{*}\rightarrow D^{*}$ be a homotopy equivalence of cochain complexes of finitely generated, free, based, left $\mathbb{Z}G$-modules. It can be regarded as a homotopy equivalence of chain complexes $f\colon C^{-*}\rightarrow D^{-*}$, and we define $\tau(f\colon C^{*}\rightarrow D^{*})\coloneqq\tau(f\colon C^{-*}\rightarrow D^{-*})$.
@@
$\operatorname{Hom}_{R}^{lr}(X,Y)$ =$=\left\{\varphi\in\operatorname{Hom}_{\mathbb{Z}}(X,Y)\mid\forall r\in R,x\in X\mid\varphi(rx)=\varphi(x)\overline{r}\right\}.$
If $Y$ has a left $S$-module structure for some ring $S$ (which is compatible with its right $R$-module structure), then $Y\otimes_{R}X$ is a left $S$-module with $s(y\otimes x)=(sy)\otimes x$, and $\operatorname{Hom}_{R}^{lr}(X,Y)$ is a left $S$-module with $(s\varphi)(x)=s\varphi(x)$.
@@
$\operatorname{Hom}_{R}^{lr}(X,Y)$ is equal to $\operatorname{Hom}_{R}^{r}(\overline{X},Y)$, the group of right $R$-module homomorphisms $\overline{X}\rightarrow Y$, as a subgroup of $\operatorname{Hom}_{\mathbb{Z}}(X,Y)$ (or as a left $S$-module). Here $\overline{X}$ denotes the right $R$-module that is equal to $X$ as an abelian group, with its multiplication given by $xr=\overline{r}\cdot x$, where $\cdot$ denotes multiplication in $X$.
@@
Suppose that $R$ is a ring with involution. Let $X$ be a left $R$-module. Its dual is the left $R$-module $X^{*}=\operatorname{Hom}_{R}^{lr}(X,R)$.
Now suppose that $G$ is equipped with a group homomorphism $w\colon G\to\{\pm 1\}$, i.e. an orientation character. This determines an involution on the group ring $\mathbb{Z}G$, and also on the Whitehead group $\operatorname{Wh}(G,w)$ (see Section 2.2). So if $C_{*}$ is a finitely generated, free, based, left $\mathbb{Z}G$-module chain complex then, using Definition 2.23, we can define the dual cochain complex $C^{*}$, which also consists of finitely generated, free, based, left $\mathbb{Z}G$-modules.
@@
Let $f\colon C_{*}\rightarrow D_{*}$ be a chain homotopy equivalence and let $f^{*}\colon D^{*}\rightarrow C^{*}$ be its dual. Then $\tau(f^{*})=\overline{\tau(f)}\in\operatorname{Wh}(G,w)$.
@@
First note that if $g\colon X\rightarrow Y$ is an isomorphism between finitely generated, free, based, left $\mathbb{Z}G$-modules, and $g^{*}\colon Y^{*}\rightarrow X^{*}$ is its dual, then the matrix of $g^{*}$ is the conjugate transpose of the matrix of $g$, hence $\tau(g^{*})=\overline{\tau(g)}\in\operatorname{Wh}(G,w)$.
Now let $(\mathscr{C}(f),\partial^{f}_{*})$ denote the mapping cone of $f$, so that $\mathscr{C}(f)_{i}=C_{i-1}\oplus D_{i}$. We regard $f^{*}\colon D^{*}\rightarrow C^{*}$ as a homotopy equivalence $f^{*}\colon D^{-*}\rightarrow C^{-*}$ of chain complexes, and define its mapping cone $(\mathscr{C}(f^{*}),\partial^{f^{*}}_{*})$. Then $\mathscr{C}(f^{*})_{i}=(D_{-i+1})^{*}\oplus(C_{-i})^{*}\cong(\mathscr{C}(f)_{-i+1})^{*}$, and this implies that $\mathscr{C}(f^{*})_{\operatorname{odd}}\cong(\mathscr{C}(f)_{\operatorname{even}})^{*}$ and $\mathscr{C}(f^{*})_{\operatorname{even}}\cong(\mathscr{C}(f)_{\operatorname{odd}})^{*}$. Moreover, $\partial^{f^{*}}_{i}\colon\mathscr{C}(f^{*})_{i}\rightarrow\mathscr{C}(f^{*})_{i-1}$ is the dual of $\partial^{f}_{-i+2}\colon\mathscr{C}(f)_{-i+2}\rightarrow\mathscr{C}(f)_{-i+1}$.
Let $s_{*}\colon\mathscr{C}(f)_{*}\rightarrow\mathscr{C}(f)_{*+1}$ be a chain contraction. We define $s^{\prime}_{*}\colon\mathscr{C}(f^{*})_{*}\rightarrow\mathscr{C}(f^{*})_{*+1}$ by $s^{\prime}_{i}=(s_{-i})^{*}\colon\mathscr{C}(f^{*})_{i}\cong(\mathscr{C}(f)_{-i+1})^{*}\rightarrow\mathscr{C}(f^{*})_{i+1}\cong(\mathscr{C}(f)_{-i})^{*}$. Then it follows by dualising the formula $s_{*}\partial^{f}+\partial^{f}s_{*}=\operatorname{Id}$ that $s^{\prime}_{*}$ is a chain contraction of $\mathscr{C}(f^{*})$.
Thus the map $\bigoplus_{i}\bigl{(}\partial^{f^{*}}_{2i+1}+s^{\prime}_{2i+1}\bigr{)}\colon\mathscr{C}(f^{*})_{\operatorname{odd}}\to\mathscr{C}(f^{*})_{\operatorname{even}}$ is the dual of the map $\bigoplus_{i}\bigl{(}\partial^{f}_{2i+1}+s_{2i+1}\bigr{)}\colon\mathscr{C}(f)_{\operatorname{odd}}\to\mathscr{C}(f)_{\operatorname{even}}$. By our earlier observations in the first paragraph of the proof, this implies that $\tau(f^{*})=\overline{\tau(f)}$. ∎
@@
    $\operatorname{Hom}_{\mathbb{Z}B}^{lr}(X_{\theta^{-1}},\mathbb{Z}B)=\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}B^{\theta})\cong\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}A)_{\theta^{-1}}$, in particular $(X_{\theta^{-1}})^{*}\cong(X^{*})_{\theta^{-1}}$.
@@
(b) (cf. [Ni23, Proposition 3.9]) For $Y$ as above, it also follows from the definitions that $\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,Y)=\operatorname{Hom}_{\mathbb{Z}B}^{lr}(X_{\theta^{-1}},Y^{\theta^{-1}})$, so $\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}B^{\theta})=\operatorname{Hom}_{\mathbb{Z}B}^{lr}(X_{\theta^{-1}},\mathbb{Z}B)=(X_{\theta^{-1}})^{*}$.
The isomorphism $\theta$ induces an isomorphism $\mathbb{Z}\theta\colon\mathbb{Z}A\rightarrow\mathbb{Z}B$ of abelian groups (or rings). This map is also an isomorphism $\mathbb{Z}\theta\colon\mathbb{Z}A\rightarrow\mathbb{Z}B^{\theta}$ of right $\mathbb{Z}A$-modules. Hence it induces an isomorphism $(\mathbb{Z}\theta)_{*}\colon\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}A)\rightarrow\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}B^{\theta})$ of abelian groups. We can check that $(\mathbb{Z}\theta)_{*}$ is also an isomorphism $\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}A)_{\theta^{-1}}\rightarrow\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}B^{\theta})$ of left $\mathbb{Z}B$-modules. So we get that $(X^{*})_{\theta^{-1}}=\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}A)_{\theta^{-1}}\cong\operatorname{Hom}_{\mathbb{Z}A}^{lr}(X,\mathbb{Z}B^{\theta})$. ∎
@@
First note that if $g\colon X\rightarrow Y$ is an isomorphism between finitely generated, free, based, left $\mathbb{Z}B$-modules, and we regard it as an isomorphism $g\colon X_{\theta}\rightarrow Y_{\theta}$ of $\mathbb{Z}A$-modules, then its matrix changes by applying $\theta^{-1}$ to each entry, hence $\tau(g\colon X_{\theta}\rightarrow Y_{\theta})=\theta^{-1}_{*}(\tau(g\colon X\rightarrow Y))\in\operatorname{Wh}(A)$.
Now let $(\mathscr{C}(f),\partial_{*})$ denote the mapping cone of $f$ and let $s_{*}\colon\mathscr{C}(f)_{*}\rightarrow\mathscr{C}(f)_{*+1}$ be the chain contraction used to define $\tau(f)$ over $\mathbb{Z}B$. Then $s_{*}$ is also a chain contraction for $(\mathscr{C}(f)_{\theta},\partial_{*})$, the mapping cone over $\mathbb{Z}A$.
So we can compute $\tau(f\colon(C_{*})_{\theta}\rightarrow(D_{*})_{\theta})$ from

### v3

@@ 2.2. The Whitehead group
Define the Whitehead group of a group $G$ to be $\operatorname{Wh}(G)=K_{1}(\mathbb{Z}G)/{\pm G}$, where the map $\pm G\to K_{1}(\mathbb{Z}G)$ is the composition $\pm G\subseteq\operatorname{GL}_{1}(\mathbb{Z}G)\subseteq\operatorname{GL}(\mathbb{Z}G)\to K_{1}(\mathbb{Z}G)$.
Note that $\operatorname{Wh}$ is a functor from the category of groups to the category of abelian groups. The map induced by a homomorphism $\theta\colon G\rightarrow H$ is denoted by $\theta_{*}\colon\operatorname{Wh}(G)\rightarrow\operatorname{Wh}(H)$. Similarly, for a continuous map $f\colon X\rightarrow Y$ between topological spaces, the map induced by $\pi_{1}(f)\colon\pi_{1}(X)\rightarrow\pi_{1}(Y)$ is denoted by $f_{*}\colon\operatorname{Wh}(\pi_{1}(X))\rightarrow\operatorname{Wh}(\pi_{1}(Y))$. The isomorphism type of $\operatorname{Wh}(G)$ is known in some cases. For us the following examples are relevant.
@@
The Whitehead group $\operatorname{Wh}(G)$ of a group $G$ is equipped with a natural involution, i.e. an automorphism $x\mapsto\overline{x}$ such that $\overline{\overline{x}}=x$. Equivalently, it has a $\mathbb{Z}C_{2}$-module structure, where the generator of $C_{2}$ acts by the involution. We describe this involution next.
@@
If $G$ is a group and $w\colon G\to\{\pm 1\}$ is an orientation character, i.e. a homomorphism from $G$ to $\{\pm 1\}$, then the integral group ring $\mathbb{Z}G$ has an involution given by $\sum_{i=1}^{k}n_{i}g_{i}\mapsto\sum_{i=1}^{k}w(g_{i})n_{i}g_{i}^{-1}$ for $n_{i}\in\mathbb{Z}$ and $g_{i}\in G$. The resulting involution on $K_{1}(\mathbb{Z}G)$ preserves $\pm G$ and so induces an involution on $\operatorname{Wh}(G)$.
We write $\operatorname{Wh}(G,w)$ for the abelian group $\operatorname{Wh}(G)$ equipped with the involution determined by the orientation character $w\colon G\to\{\pm 1\}$. This can equivalently be regarded as a $\mathbb{Z}C_{2}$-module. When $w\equiv 1$ is the trivial homomorphism, we omit it from the notation and write $\operatorname{Wh}(G)$ for $\operatorname{Wh}(G,1)$. Note that the choice of $w$ only affects the involution, and $\operatorname{Wh}(G,w)$ is equal to $\operatorname{Wh}(G)$ as an abelian group.
Next we recall, from the introduction, the definition of two subgroups of $\operatorname{Wh}(G,w)$ which will play an important rôle in the rest of this article.
@@
For a group $G$, a homomorphism $w\colon G\to\{\pm 1\}$ and $n\in\mathbb{Z}$, define:
$\mathcal{J}_{n}(G,w)$ $=\{y\in\operatorname{Wh}(G,w)\mid y=-(-1)^{n}\overline{y}\}\leq\operatorname{Wh}(G,w)$
$\mathcal{I}_{n}(G,w)$ $=\{x-(-1)^{n}\overline{x}\mid x\in\operatorname{Wh}(G,w)\}\leq\operatorname{Wh}(G,w).$
We have $\mathcal{I}_{n}(G,w)\leq\mathcal{J}_{n}(G,w)$, and it follows from Proposition 9.4 that there is a canonical isomorphism
@@ 2.3. The Whitehead torsion of a chain homotopy equivalence
We now define the Whitehead torsion in the algebraic setting. A free, based $R$-module is a free $R$-module $P$ together with a choice of isomorphism $P\cong R^{k}$ for some $k$. The canonical basis for $R^{k}$ determines a basis for $P$.
Let $G$ be a group and let $f\colon C_{*}\to D_{*}$ be a chain homotopy equivalence between finitely generated, free, based (left) $\mathbb{Z}G$-module chain complexes. Consider the algebraic mapping cone $(\mathscr{C}(f),\partial_{*})$, which is also a finitely generated, free, based $\mathbb{Z}G$-module chain complex. Since $f$ is a chain homotopy equivalence, $\mathscr{C}(f)$ is chain contractible; see e.g. [Ranicki-AGS, Proposition 3.14]. Choose a chain contraction $s\colon\mathscr{C}(f)_{*}\to\mathscr{C}(f)_{*+1}$. Now consider the modules
$\mathscr{C}(f)_{\operatorname{odd}}:=\bigoplus_{i=0}^{\infty}\mathscr{C}(f)_{2i+1}\quad\text{ and }\quad\mathscr{C}(f)_{\operatorname{even}}:=\bigoplus_{i=0}^{\infty}\mathscr{C}(f)_{2i}.$
@@
$\tau(f):=\Bigl[\bigoplus_{i=0}^{\infty}\bigl(\partial_{2i+1}+s_{2i+1}\bigr)\colon\mathscr{C}(f)_{\operatorname{odd}}\to\mathscr{C}(f)_{\operatorname{even}}\Bigr]\in\operatorname{Wh}(G).$
This equivalence class is independent of the choice of the chain contraction $s$ [Co73, (15.3)]. If $\tau(f)=0$ then we say that $f$ is a simple chain homotopy equivalence.
@@
Since $\mathscr{C}(f)_{\operatorname{odd}}$ and $\mathscr{C}(f)_{\operatorname{even}}$ are finitely generated, the sum in the definition of $\tau(f)$ is a finite sum. The equivalence class $\tau(f)$ remains invariant under permutations of the bases of $C_{*}$ and $D_{*}$, or if a basis element is multiplied by (1)(-1) or an element of $G$ [Co73, (15.2) and (10.3)].
We now list some useful facts about $\tau$. We fix the group $G$, and assume all chain complexes are finitely generated, free, based, left $\mathbb{Z}G$-module chain complexes.
@@
For a chain complex $C_{*}$ and an integer $k$ we denote by $C_{k+*}$ the same chain complex with shifted grading $i\mapsto C_{k+i}$. Similarly, the cochain complex $C_{k-*}$ is defined by changing the grading to $i\mapsto C_{k-i}$. The notation $C^{k+i}$ and $C^{k-*}$ is defined analogously for a cochain complex $C^{*}$.
@@
Let $f\colon C^{*}\rightarrow D^{*}$ be a homotopy equivalence of cochain complexes of finitely generated, free, based, left $\mathbb{Z}G$-modules. It can be regarded as a homotopy equivalence of chain complexes $f\colon C^{-*}\rightarrow D^{-*}$, and we define $\tau(f\colon C^{*}\rightarrow D^{*}):=\tau(f\colon C^{-*}\rightarrow D^{-*})$.
@@
$\operatorname{Hom}_{R}^{\mathrm{lr}}(X,Y)$ $=\left\{\varphi\in\operatorname{Hom}_{\mathbb{Z}}(X,Y)\mid\forall r\in R,x\in X\colon\varphi(rx)=\varphi(x)\overline{r}\right\}.$
If $Y$ is moreover an $(S,R)$-bimodule for some ring $S$, then $Y\otimes_{R}X$ is a left $S$-module with $s(y\otimes x)=(sy)\otimes x$, and $\operatorname{Hom}_{R}^{\mathrm{lr}}(X,Y)$ is a left $S$-module with $(s\varphi)(x)=s\varphi(x)$.
@@
$\operatorname{Hom}_{R}^{\mathrm{lr}}(X,Y)$ is equal to $\operatorname{Hom}_{R}^{r}(\overline{X},Y)$, the group of right $R$-module homomorphisms $\overline{X}\rightarrow Y$, as a subgroup of $\operatorname{Hom}_{\mathbb{Z}}(X,Y)$ (or as a left $S$-module). Here $\overline{X}$ denotes the right $R$-module that is equal to $X$ as an abelian group, with its multiplication given by $xr=\overline{r}\cdot x$, where $\cdot$ denotes multiplication in $X$.
@@
Suppose that $R$ is a ring with involution. Let $X$ be a left $R$-module. Its dual is the left $R$-module $X^{*}:=\operatorname{Hom}_{R}^{\mathrm{lr}}(X,R)$.
Now suppose that $G$ is equipped with a group homomorphism $w\colon G\to\{\pm 1\}$, i.e. an orientation character. This determines an involution on the group ring $\mathbb{Z}G$, and also on the Whitehead group $\operatorname{Wh}(G,w)$ (see Section 2.2). Since $\operatorname{Hom}_{R}^{\mathrm{lr}}$ depends on the involution of $R$, we will write $\operatorname{Hom}_{\mathbb{Z}G,w}^{\mathrm{lr}}$ when $\mathbb{Z}G$ is equipped with the involution determined by $w$, and $\operatorname{Hom}_{\mathbb{Z}G}^{\mathrm{lr}}$ will mean $\operatorname{Hom}_{\mathbb{Z}G,1}^{\mathrm{lr}}$.
Definition 2.21.
We introduce the following notation.
  • $\bullet$
    Let $\mathbb{Z}^{w}$ be the right $\mathbb{Z}G$-module with underlying abelian group $\mathbb{Z}$ and multiplication $ng=w(g)n$, where $n\in\mathbb{Z}$, $g\in G$ and $w(g)\in\{\pm 1\}\subseteq\mathbb{Z}$.
  • $\bullet$
    For a right $\mathbb{Z}G$-module $Y$ we define the right $\mathbb{Z}G$-module $Y^{w}:=\mathbb{Z}^{w}\otimes_{\mathbb{Z}}Y$, with underlying abelian group $\mathbb{Z}\otimes_{\mathbb{Z}}Y$ and multiplication $(n\otimes y)g=(ng)\otimes(yg)$, where $n\in\mathbb{Z}$, $y\in Y$ and $g\in G$. Equivalently, $Y^{w}$ has underlying abelian group $Y$ and multiplication $yg=w(g)y\cdot g$, where $\cdot$ is the multiplication in $Y$. If $Y$ is also a left $S$-module for some ring $S$, then so is $Y^{w}$.
For example, $\mathbb{Z}G^{w}$ is a $\mathbb{Z}G$-bimodule, whose left multiplication is the same as that of $\mathbb{Z}G$, and whose right multiplication is twisted by $w$.
If $X$ is a left $\mathbb{Z}G$-module, then by the definitions we have $\operatorname{Hom}_{\mathbb{Z}G,w}^{\mathrm{lr}}(X,Y)=\operatorname{Hom}_{\mathbb{Z}G}^{\mathrm{lr}}(X,Y^{w})$ as abelian groups (or as left $S$-modules, when $Y$ is a left $S$-module).
Definition 2.22.
For a group $G$, homomorphism $w\colon G\to\{\pm 1\}$ and left $\mathbb{Z}G$-module $X$, the left $\mathbb{Z}G$-module $\operatorname{Hom}_{\mathbb{Z}G,w}^{\mathrm{lr}}(X,\mathbb{Z}G)=\operatorname{Hom}_{\mathbb{Z}G}^{\mathrm{lr}}(X,\mathbb{Z}G^{w})$ will be denoted by $X^{w*}$.
If $X$ is a finitely generated, free, based, left $\mathbb{Z}G$-module, then $X^{w*}$ is also a finitely generated, free, left $\mathbb{Z}G$-module, naturally equipped with a dual basis. If $f\colon X_{1}\rightarrow X_{2}$ is a homomorphism between two such modules, then the matrix of its dual, $f^{w*}\colon X_{2}^{w*}\rightarrow X_{1}^{w*}$ is the conjugate transpose of the matrix of $f$, where conjugation means applying the involution determined by $w$.
If $C_{*}$ is a finitely generated, free, based, left $\mathbb{Z}G$-module chain complex, then $C^{w*}$ will denote the dual cochain complex, which also consists of finitely generated, free, based, left $\mathbb{Z}G$-modules.
@@
Let $f\colon C_{*}\rightarrow D_{*}$ be a chain homotopy equivalence and let $f^{w*}\colon D^{w*}\rightarrow C^{w*}$ be its dual. Then $\tau(f^{w*})=\overline{\tau(f)}\in\operatorname{Wh}(G,w)$.
@@
Let $(\mathscr{C}(f),\partial^{f}_{*})$ denote the mapping cone of $f$, so that $\mathscr{C}(f)_{i}=C_{i-1}\oplus D_{i}$. We regard $f^{w*}\colon D^{w*}\rightarrow C^{w*}$ as a homotopy equivalence $f^{w*}\colon D^{w(-*)}\rightarrow C^{w(-*)}$ of chain complexes, and define its mapping cone $(\mathscr{C}(f^{w*}),\partial^{f^{w*}}_{*})$. Then
$\mathscr{C}(f^{w*})_{i}=(D_{-i+1})^{w*}\oplus(C_{-i})^{w*}\cong(\mathscr{C}(f)_{-i+1})^{w*},$
and this implies that $\mathscr{C}(f^{w*})_{\operatorname{odd}}\cong(\mathscr{C}(f)_{\operatorname{even}})^{w*}$ and $\mathscr{C}(f^{w*})_{\operatorname{even}}\cong(\mathscr{C}(f)_{\operatorname{odd}})^{w*}$. Moreover, $\partial^{f^{w*}}_{i}\colon\mathscr{C}(f^{w*})_{i}\rightarrow\mathscr{C}(f^{w*})_{i-1}$ is the dual of $\partial^{f}_{-i+2}\colon\mathscr{C}(f)_{-i+2}\rightarrow\mathscr{C}(f)_{-i+1}$.
Let $s_{*}\colon\mathscr{C}(f)_{*}\rightarrow\mathscr{C}(f)_{*+1}$ be a chain contraction. We define $s^{\prime}_{*}\colon\mathscr{C}(f^{w*})_{*}\rightarrow\mathscr{C}(f^{w*})_{*+1}$ by
$s^{\prime}_{i}=(s_{-i})^{w*}\colon\mathscr{C}(f^{w*})_{i}\cong(\mathscr{C}(f)_{-i+1})^{w*}\rightarrow\mathscr{C}(f^{w*})_{i+1}\cong(\mathscr{C}(f)_{-i})^{w*}.$
Then it follows by dualising the formula $s_{*}\partial^{f}+\partial^{f}s_{*}=\operatorname{Id}$ that $s^{\prime}_{*}$ is a chain contraction of $\mathscr{C}(f^{w*})$. Thus $\bigoplus_{i}\bigl(\partial^{f^{w*}}_{2i+1}+s^{\prime}_{2i+1}\bigr)\colon\mathscr{C}(f^{w*})_{\operatorname{odd}}\to\mathscr{C}(f^{w*})_{\operatorname{even}}$ is the dual of $\bigoplus_{i}\bigl(\partial^{f}_{2i+1}+s_{2i+1}\bigr)\colon\mathscr{C}(f)_{\operatorname{odd}}\to\mathscr{C}(f)_{\operatorname{even}}$. Hence its matrix is the conjugate transpose of the matrix of the latter map, which implies that $\tau(f^{w*})=\overline{\tau(f)}$. ∎
@@
    $\operatorname{Hom}_{\mathbb{Z}B,w_{B}}^{\mathrm{lr}}(X_{\theta^{-1}},\mathbb{Z}B)=\operatorname{Hom}_{\mathbb{Z}A,w_{A}}^{\mathrm{lr}}(X,\mathbb{Z}B^{\theta})\cong\operatorname{Hom}_{\mathbb{Z}A,w_{A}}^{\mathrm{lr}}(X,\mathbb{Z}A)_{\theta^{-1}}$, in particular we have $(X_{\theta^{-1}})^{w_{B}*}\cong(X^{w_{A}*})_{\theta^{-1}}$.
@@
(b) This follows from [Ni23, Proposition 3.9]. ∎
@@
First note that if $g\colon X\rightarrow Y$ is an isomorphism between finitely generated, free, based, left $\mathbb{Z}B$-modules, and we regard it as an isomorphism $g\colon X_{\theta}\rightarrow Y_{\theta}$ of $\mathbb{Z}A$-modules, then its matrix changes by applying $\mathbb{Z}\theta^{-1}$ to each entry, hence $\tau(g\colon X_{\theta}\rightarrow Y_{\theta})=\theta^{-1}_{*}(\tau(g\colon X\rightarrow Y))\in\operatorname{Wh}(A)$.
Now let $(\mathscr{C}(f),\partial_{*})$ denote the mapping cone of $f$ and let $s_{*}\colon\mathscr{C}(f)_{*}\rightarrow\mathscr{C}(f)_{*+1}$ be the chain contraction used to define $\tau(f)$ over $\mathbb{Z}B$. Then $s_{*}$ is also a chain contraction for $(\mathscr{C}(f)_{\theta},\partial_{*})$, the mapping cone over $\mathbb{Z}A$. So we can compute $\tau(f\colon(C_{*})_{\theta}\rightarrow(D_{*})_{\theta})$ from